# Task-Agnostic Machine Learning-Assisted Inference

### v1

@@ Abstract
Machine learning (ML) is playing an increasingly important role in scientific research. In conjunction with classical statistical approaches, ML-assisted analytical strategies have shown great promise in accelerating research findings. This has also opened up a whole new field of methodological research focusing on integrative approaches that leverage both ML and statistics to tackle data science challenges. One type of study that has quickly gained popularity employs ML to predict unobserved outcomes in massive samples and then uses the predicted outcomes in downstream statistical inference. However, existing methods designed to ensure the validity of this type of post-prediction inference are limited to very basic tasks such as linear regression analysis. This is because any extension of these approaches to new, more sophisticated statistical tasks requires task-specific algebraic derivations and software implementations, which ignores the massive library of existing software tools already developed for complex inference tasks and severely constrains the scope of post-prediction inference in real applications. To address this challenge, we propose a novel statistical framework for task-agnostic ML-assisted inference. It provides a post-prediction inference solution that can be easily plugged into almost any established data analysis routine. It delivers valid and efficient inference that is robust to arbitrary choices of ML models, while allowing nearly all existing analytical frameworks to be incorporated into the analysis of ML-predicted outcomes. Through extensive experiments, we showcase the validity, versatility, and superiority of our method compared to existing approaches.
@@ 1 Introduction
Leveraging machine learning (ML) techniques to enhance and accelerate research has become increasingly popular in many scientific disciplines (Wang et al.,, 2023). For example, sophisticated deep learning frameworks have achieved remarkable success in predicting protein structure and interactions, which has the potential to significantly speed up the research process, save costs, and revolutionize the field of structural biology (Jumper et al.,, 2021; Abramson et al.,, 2024; Ahdritz et al.,, 2024). However, several recent studies have pointed out that statistical inference using ML-predicted outcomes may lead to invalid scientific discoveries due to the lack of consideration of ML prediction uncertainty in traditional statistical approaches (Wang et al.,, 2020; Angelopoulos et al., 2023a, ; Miao et al., 2023b, ). To address this, researchers have introduced methods that couple ML prediction on large samples and a small set of individuals with measured outcome data to ensure the validity of ML-assisted statistical inference (Wang et al.,, 2020; Angelopoulos et al., 2023a, ; Miao et al., 2023b, ).
@@
Despite these advances, current ML-assisted inference methods can only address basic statistical tasks, including mean estimation, quantile estimation, and linear and logistic regression (Angelopoulos et al., 2023a, ; Miao et al., 2023b, ). While the same mathematical principle behind existing ML-assisted inference methods can be generalized to a broader class of M-estimation problems, specific algebraic derivations and computational implementations are required for each new statistical task. Moreover, many tasks, such as the Wilcoxon rank-sum test, do not fit into the M-estimation framework. These issues pose significant challenges to the broad application of ML-assisted inference across various scientific domains.
@@
Historically, the statistics field has encountered similar challenges before the advent of resampling-based methods (Efron and Tibshirani,, 1994). It used to require task-specific derivation and implementation in order to obtain the standard error of any new estimator. This process is remarkably similar to the current state of ML-assisted inference, where every new task requires researchers to go through these non-trivial steps. However, with resampling-based inference, the need to manually derive standard errors was reduced. Instead, resampling methods can be universally applied to many estimation problems and easily obtain standard errors (Efron,, 1982, 1992; Efron and Tibshirani,, 1994). Inspired by this, we seek a universal approach that can incorporate ML-predicted outcomes into any existing analysis routine while ensuring the validity of inference results.
@@
We introduce a simple protocol named PoSt-Prediction Sumstats-based (PSPS) inference. It employs existing analysis routines to generate summary statistics that are sufficient for ML-assisted inference, then produces valid inference results using these summary statistics. We summarize the key features of our method:
@@
  • Assumption-lean and data-adaptive: It inherits the theoretical guarantees, including validity and efficiency, of the state-of-the-art methods for ML-assisted inference (Angelopoulos et al., 2023b, ; Miao et al., 2023b, ; Gan and Liang,, 2023). These guarantees are robust to the choice of ML procedure.
@@
  • Task-agnostic and simple: Due to the fact that our method only needs summary statistics produced by existing analysis routines as its input, it can be easily adopted for a broad collection of statistical tasks that are currently unavailable or would require substantial effort to implement in the ML-assisted inference setting.
@@
  • Privacy-preserving: It does not need any individual-level data as input. Sharing of privacy-preserving summary statistics is sufficient for real-world scientific collaboration.
The rest of this paper is organized as follows. section 2 formulates the ML-assisted inference problem and builds the intuition for our method. section 3 develops our general protocol for PSPS, establishes its asymptotic properties and optimality, and provides an extension to handle different data structures. section 4 introduces ML-assisted false discovery rate control for multiple testing correction. We then evaluate our methods on simulated and real datasets in section 5 and conclude in section 6. All proofs are deferred to appendix A.
@@ 2.1 Setting
We focus on statistical inference problems for the parameter $\theta^{*}\equiv\theta^{*}(\mathbb{P})\in\mathbb{R}^{K}$ defined on the the joint distribution of $({\bf X},Y)\sim\mathbb{P}$, where $Y\in\mathcal{Y}$ is a scalar outcome and ${\bf X}\in\mathcal{X}$ be a $K$-dimensional vector representing features. We are interested in estimating $\theta^{*}$ using labeled data =${\cal L}=\{({\bf X}_{i},Y_{i}),i=1,\cdots,n\}\equiv({\bf X}_{\cal L},Y_{\cal L})$, unlabeled data ${\cal U}=\{{\bf X}_{i},i=n+1,\cdots,n+N\}\equiv{\bf X}_{\cal U}$, and a pre-trained ML model $\widehat{f}(\cdot):\mathcal{X}\rightarrow\mathcal{Y}$. Here, $f(\cdot)$ is a black-box function with unknown operating characteristics and can be mis-specified. We also require an algorithm $\mathcal{A}$ that inputs the labeled data ${\cal L}$ and returns a consistent and asymptotically normally distributed estimator $\widehat{\theta}$ for $\theta^{*}$. There are three common ways in the literature to estimate $\theta^{*}$.
  • Classical statistical methods apply algorithm $\mathcal{A}$ to only labeled data ${\cal L}=({\bf X}_{\cal L},Y_{\cal L})$, and returns the estimator and its corresponding standard error $[\widehat{\bm{\theta}}_{\cal L},\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L})]$. Valid confidence intervals and hypothesis tests can then be constructed using the asymptotic distribution of the estimator. However, it ignores the unlabeled data and ML predictions.
  • Imputation-based methods that treat the ML prediction $\widehat{f}$ in the unlabeled data as the observed outcome, and apply algorithm $\mathcal{A}$ to ${\cal U}=({\bf X}_{\cal U},\widehat{f}_{\cal U})$. We denote the estimator and corresponding standard error as $[\widehat{\bm{\eta}}_{\cal U},\operatorname{Var}(\widehat{\bm{\eta}}_{\cal U})]$. This has been shown to give invalid inference results and false scientific findings (Wang et al.,, 2020; Angelopoulos et al., 2023a, ; Miao et al., 2023b, ; Miao et al.,, 2024).
  • ML-assisted inference methods use both ${\cal L}$ and ${\cal U}$ as input. These approaches add a debiasing term in the loss function (or estimating equation) for M-estimation problems, thus removing biases from the imputation-based estimators and producing results that are statistically valid and universally more powerful compared to classical methods.
We first provide some intuition about ML-assisted inference and our proposed protocol using an example.
@@ 2.2 Building the intuition with mean estimation
We consider the mean estimation problem where $\theta^{*}=\mathbb{E}[Y_{i}]\equiv\mbox{arg min}_{\theta}\mathbb{E}[\frac{1}{2}\left(Y_{i}-\theta\right)^{2}]$. The classical method only takes the labeled data $Y_{\cal L}$ as input and yields an unbiased and consistent estimator for $\theta^{*}$:
$\widehat{\theta}_{\cal L}=\mbox{arg min}_{\theta}\frac{1}{n}\sum_{i=1}^{n}\frac{1}{2}\left(Y_{i}-\theta\right)^{2}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}.$ (1)
The imputation-based method only takes the unlabeled data $\widehat{f}_{\cal U}$ as input and returns
$\widehat{\eta}_{\cal U}=\mbox{arg min}_{\theta}\frac{1}{N}\sum_{i=n+1}^{n+N}\frac{1}{2}\left(\widehat{f}_{i}-\theta\right)^{2}=\frac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}_{i}.$ (2)
@@
It is a biased and inconsistent estimator for $\mathbb{E}[Y_{i}]$ if the ML model $\widehat{f}$ is mis-specified. To address this, the ML-assisted estimator takes both labeled data $(Y_{\cal L},\widehat{f}_{\cal L})$ and unlabeled data $\widehat{f}_{\cal U}$ as input and adds a debiasing term to the loss function to rectify the bias caused by ML imputation:
@@
$\widehat{\theta}_{\texttt{MLA}}$ $=\mbox{arg min}_{\theta}\frac{1}{2}\left\{\omega_{0}\frac{1}{N}\sum_{i=n+1}^{n+N}\left(\widehat{f}_{i}-\theta\right)^{2}-\underbrace{\left[\omega_{0}\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}_{i}-\theta\right)^{2}-\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}-\theta\right)^{2}\right]}_{\text{Debiasing term}}\right\}$ (3)
@@
$=\omega_{0}\frac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}_{i}-\underbrace{\left[\omega_{0}\frac{1}{n}\sum_{i=1}^{n}\widehat{f}_{i}-\frac{1}{n}\sum_{i=1}^{n}y_{i}\right]}_{\text{Debiasing term}},$ (4)
@@
where the added debiased loss ensures the consistency of the ML-assisted estimator, and the weight $\omega_{0}=\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]/n}{\textnormal{Var}_{l}[\widehat{f}]/n+\textnormal{Var}_{u}[\widehat{f}]/N}$ ensures that the ML-assisted estimator is no less efficient than the classical estimator with arbitrary ML predictions. $\operatorname{Var}(\widehat{\theta}_{\texttt{MLA}})=\operatorname{Var}(\widehat{\theta}_{\cal L})-\frac{\operatorname{Cov}_{l}[Y,\widehat{f}]}{n\operatorname{Var}_{l}[\widehat{f}]+n^{2}\operatorname{Var}_{u}[\widehat{f}]/N}\leq\operatorname{Var}(\widehat{\theta}_{\cal L})$.
@@
Our proposed method is motivated by the observation that the sufficient statistics of both the ML-assisted estimator $\widehat{\theta}_{\texttt{MLA}}$ and its variance $\operatorname{Var}(\widehat{\theta}_{\texttt{MLA}})$ are the following summary statistics:
@@
$\widehat{\bm{\theta}}_{\texttt{ss}}=(\widehat{\theta}_{{\cal L}},\widehat{\eta}_{{\cal L}},\widehat{\eta}_{{\cal U}})\equiv(\frac{1}{n}\sum_{i=1}^{n}y_{i},\frac{1}{n}\sum_{i=1}^{n}\widehat{f}_{i},\frac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}_{i})\text{ and }\bm{\Sigma}_{\texttt{ss}}=\operatorname{Var}(\widehat{\bm{\theta}}_{\texttt{ss}})\equiv\begin{bmatrix}\textnormal{Var}_{l}[Y]/n&\textnormal{Cov}_{l}[Y,\widehat{f}]/n&0\\ \textnormal{Cov}_{l}[Y,\widehat{f}]/n&\textnormal{Var}_{l}[\widehat{f}]/n&0\\ 0&0&\textnormal{Var}_{u}[\widehat{f}]/N\\ \end{bmatrix}$ (5)
@@
Moreover, they can be easily obtained by applying the same algorithm $\mathcal{A}$ (i.e., mean estimation in this example) to:
@@
  • labeled data with observed outcome $\mathcal{A}(Y_{\cal L})\rightarrow[\widehat{\theta}_{{\cal L}},\operatorname{Var}(\widehat{\theta}_{{\cal L}})]=[\frac{1}{n}\sum_{i=1}^{n}y_{i},\textnormal{Var}_{l}[Y]/n]$
@@
  • labeled data with predicted outcome $\mathcal{A}(\widehat{f}_{\cal L})\rightarrow[\widehat{\eta}_{{\cal L}},\operatorname{Var}(\widehat{\eta}_{{\cal L}})]=[\frac{1}{n}\sum_{i=1}^{n}\widehat{f}_{i},\textnormal{Var}_{l}[\widehat{f}]/n]$
  • unlabeled data with predicted outcome $\mathcal{A}(\widehat{f}_{\cal U})\rightarrow[\widehat{\eta}_{{\cal U}},\operatorname{Var}(\widehat{\eta}_{{\cal U}})]=[\frac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}_{i},\textnormal{Var}_{u}[\widehat{f}]/N]$
@@
  • bootstrap samples of labeled data $\mathcal{A}[(Y_{\cal L},\widehat{f}_{\cal L})_{q},q=1,\dots,Q]$ for estimation of $\operatorname{Cov}(\widehat{\theta}_{{\cal L}},\widehat{\eta}_{{\cal L}})=\textnormal{Cov}_{l}[Y,\widehat{f}]/n$. Here, $(Y_{\cal L},\widehat{f}_{\cal L})_{q}$ represents the $q$-th bootstrap of labeled data.
@@
Inputting these summary statistics for one-step debiasing $\omega_{0}\widehat{\eta}_{{\cal U}}-(\omega_{0}\widehat{\eta}_{{\cal L}}-\widehat{\theta}_{{\cal L}})$ recovers $\widehat{\theta}_{\texttt{MLA}}$.
@@
To summarize, an algorithm for mean estimation, coupled with resampling, is sufficient for ML-assisted estimation. This inspires us to generalize this protocol for a broad range of statistical tasks. Our protocol illustrated in fig. 1 only requires three steps: 1) using the pre-trained ML model to predict the outcomes in both the labeled and unlabeled data, 2) applying the existing analysis routines to generate summary statistics, 3) employing a debiasing procedure using these summary statistics and producing statistically valid results.
Figure 1: Workflow of PSPS for Task-Agnostic ML-Assisted Inference

### v3

# Task-Agnostic Machine-Learning-Assisted Inference

@@ Abstract
Machine learning (ML) is playing an increasingly important role in scientific research. In conjunction with classical statistical approaches, ML-assisted analytical strategies have shown great promise in accelerating research findings. This has also opened a whole field of methodological research focusing on integrative approaches that leverage both ML and statistics to tackle data science challenges. One type of study that has quickly gained popularity employs ML to predict unobserved outcomes in massive samples, and then uses predicted outcomes in downstream statistical inference. However, existing methods designed to ensure the validity of this type of post-prediction inference are limited to very basic tasks such as linear regression analysis. This is because any extension of these approaches to new, more sophisticated statistical tasks requires task-specific algebraic derivations and software implementations, which ignores the massive library of existing software tools already developed for the same scientific problem given observed data. This severely constrains the scope of application for post-prediction inference. To address this challenge, we introduce a novel statistical framework named PSPS for task-agnostic ML-assisted inference. It provides a post-prediction inference solution that can be easily plugged into almost any established data analysis routines. It delivers valid and efficient inference that is robust to arbitrary choice of ML model, allowing nearly all existing statistical frameworks to be incorporated into the analysis of ML-predicted data. Through extensive experiments, we showcase our method’s validity, versatility, and superiority compared to existing approaches. Our software is available at https://github.com/qlu-lab/psps.
@@ 1 Introduction
Leveraging machine learning (ML) techniques to enhance and accelerate research has become increasingly popular in many scientific disciplines [44]. For example, sophisticated deep learning models have achieved remarkable success in predicting protein structure and interactions, which has the potential to significantly speed up the research process, save costs, and revolutionize the field of structural biology [25, 1, 2]. However, recent studies have pointed out that statistical inference using ML-predicted outcomes may lead to invalid scientific discoveries due to the lack of consideration of ML prediction uncertainty in traditional statistical approaches. To address this, researchers have introduced methods that couple extensive ML predictions with limited gold-standard data to ensure the validity of ML-assisted statistical inference [46, 3, 35].
@@
Despite these advances, current ML-assisted inference methods can only address very basic statistical tasks, including mean estimation, quantile estimation, and linear and logistic regression [3, 35]. While the same mathematical principle behind existing ML-assisted inference methods can be generalized to a broader class of M-estimation problems, specific algebraic derivations and computational implementations are required for each new statistical task. Moreover, many tasks, such as the Wilcoxon rank-sum test, do not fit into the M-estimation framework. These issues pose significant challenges to the broad application of ML-assisted inference across various scientific domains.
@@
Historically, the field of statistics has faced similar types of challenges. Before the advent of resampling-based methods [19], it used to require task-specific derivation and implementation to obtain the variance of any new estimator. This old problem mirrors the current state of ML-assisted inference, where every new task requires non-trivial effort from researchers. However, with resampling-based inference, the need to manually derive variance is reduced. Instead, resampling methods can be universally applied to many estimation problems and easily obtain variance [17, 18, 19]. Inspired by this, we seek a universal approach that incorporates ML-predicted data into any existing data analysis routines while ensuring valid inference results.
@@
We introduce a simple protocol named PoSt-Prediction Summary-statistics-based (PSPS) inference (fig. 1). It employs existing analysis routines to generate summary statistics sufficient for ML-assisted inference, and then produces valid and powerful inference results using these statistics. It has several key features:
@@
  • Assumption-lean and data-adaptive: It inherits the theoretical guarantees of validity and efficiency from state-of-the-art ML-assisted inference methods [4, 35, 20]. These guarantees hold with arbitrary ML predictions.
@@
  • Task-agnostic and simple: Since our method only requires summary statistics from existing analysis routines, it can be easily adapted for many statistical tasks currently unavailable or difficult to implement in ML-assisted inference.
@@
  • Federated data analysis: It does not need any individual-level data as input. Sharing of privacy-preserving summary statistics is sufficient for real-world scientific collaboration.
@@ 2.1 Setting
We focus on statistical inference problems for the parameter $\theta^{*}\equiv\theta^{*}(\mathbb{P})\in\mathbb{R}^{K}$ defined on the joint distribution of $({\bf X},Y)\sim\mathbb{P}$, where $Y\in\mathcal{Y}$ is a scalar outcome and ${\bf X}\in\mathcal{X}$ be a $K$-dimensional vector representing features. We are interested in estimating $\theta^{*}$ using labeled data =${\cal L}=\{({\bf X}_{i},Y_{i}),i=1,\cdots,n\}\equiv({\bf X}_{\cal L},Y_{\cal L})$, unlabeled data  ${\cal U}=\{{\bf X}_{i},i=n+1,\cdots,n+N\}\equiv{\bf X}_{\cal U}$, and a pre-trained ML model $\widehat{f}(\cdot):\mathcal{X}\rightarrow\mathcal{Y}$. Here, $f(\cdot)$ is a black-box function with unknown operating characteristics and can be mis-specified. We also require an algorithm $\mathcal{A}$ that inputs the labeled data ${\cal L}$ and returns a consistent and asymptotically normally distributed estimator $\widehat{\theta}$ for $\theta^{*}$. There are three common ways in the literature to estimate $\theta^{*}$:
  • Classical statistical methods apply algorithm $\mathcal{A}$ to only labeled data ${\cal L}=({\bf X}_{\cal L},Y_{\cal L})$, and returns the estimator and its estimated variance $[\widehat{\bm{\theta}}_{\cal L},\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\cal L})]$. Valid confidence intervals and hypothesis tests can then be constructed using the asymptotic distribution of the estimator. However, it ignores the unlabeled data and ML prediction.
  • Imputation-based methods treat ML prediction $\widehat{f}$ in the unlabeled data as the observed outcome, and apply algorithm $\mathcal{A}$ to ${\cal U}=({\bf X}_{\cal U},\widehat{f}_{\cal U})$. We denote the estimator and estimated variance as $[\widehat{\bm{\eta}}_{\cal U},\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{\cal U})]$. This has been shown to give invalid inference results and false scientific findings [46, 3, 35, 36].
  • ML-assisted inference methods use both ${\cal L}$ and ${\cal U}$ as input. These approaches add a debiasing term in the loss function (or estimating equation) for M-estimation problems, thus removing the bias from the imputation-based estimators and producing results that are statistically valid and universally more powerful compared to classical methods [4, 35, 36].
Next, we use an example to provide intuition on ML-assisted inference and our protocol.
@@ 2.2 Building the intuition with mean estimation
We consider the mean estimation problem, where $\theta^{*}=\mathbb{E}[Y_{i}]\equiv\mbox{arg min}_{\theta}\mathbb{E}[\frac{1}{2}\left(Y_{i}-\theta\right)^{2}]$. The classical method only takes the labeled data $Y_{\cal L}$ as input and yields an unbiased and consistent estimator for $\theta^{*}$: $\widehat{\theta}_{\cal L}=\mbox{arg min}_{\theta}\frac{1}{n}\sum\limits_{i=1}^{n}\frac{1}{2}\left(Y_{i}-\theta\right)^{2}=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}$. The imputation-based method only takes the unlabeled data $\widehat{f}_{\cal U}$ as input and returns $\widehat{\eta}_{\cal U}=\mbox{arg min}_{\theta}\frac{1}{N}\sum_{i=n+1}^{n+N}\frac{1}{2}(\widehat{f}_{i}-\theta)^{2}=\frac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}_{i}$. It is a biased and inconsistent estimator for $\mathbb{E}[Y_{i}]$ if the ML model $\widehat{f}$ is mis-specified. To address this, ML-assisted estimator takes both labeled data $(Y_{\cal L},\widehat{f}_{\cal L})$ and unlabeled data $\widehat{f}_{\cal U}$ as input and adds a debiasing term to the loss function to rectify the bias caused by ML imputation [3, 35, 36, 20]:
@@
$\widehat{\theta}_{\texttt{MLA}}$ $=\mbox{arg min}_{\theta}\frac{1}{2}\{\widehat{\omega}_{0}\frac{1}{N}\sum\limits_{i=n+1}^{n+N}(\widehat{f}_{i}-\theta)^{2}-\underbrace{[\widehat{\omega}_{0}\frac{1}{n}\sum\limits_{i=1}^{n}(\widehat{f}_{i}-\theta)^{2}-\frac{1}{n}\sum\limits_{i=1}^{n}\left(Y_{i}-\theta\right)^{2}]}_{\text{Debiasing term}}\}$
@@
$=\widehat{\omega}_{0}\frac{1}{N}\sum\limits_{i=n+1}^{n+N}\widehat{f}_{i}-\underbrace{[\widehat{\omega}_{0}\frac{1}{n}\sum\limits_{i=1}^{n}\widehat{f}_{i}-\frac{1}{n}\sum\limits_{i=1}^{n}y_{i}]}_{\text{Debiasing term}},$
@@
where the modified loss ensures the consistency of the ML-assisted estimator and the weight $\widehat{\omega}_{0}=\frac{\widehat{\textnormal{Cov}}_{l}[Y,\widehat{f}]/n}{\widehat{\textnormal{Var}}_{l}[\widehat{f}]/n+\widehat{\textnormal{Var}}_{u}[\widehat{f}]/N}$ ensures that ML-assisted estimator is no less efficient than the classical estimator with arbitrary ML predictions: $\operatorname{Var}(\widehat{\theta}_{\texttt{MLA}})=\operatorname{Var}(\widehat{\theta}_{\cal L})-\frac{\operatorname{Cov}[Y,\widehat{f}]}{n\operatorname{Var}[\widehat{f}]+n^{2}\operatorname{Var}[\widehat{f}]/N}\leq\operatorname{Var}(\widehat{\theta}_{\cal L})$.
@@
Our proposed method is motivated by the observation that the sufficient statistics of the ML-assisted estimator $\widehat{\theta}_{\texttt{MLA}}$ and its estimated variance $\widehat{\operatorname{Var}}(\widehat{\theta}_{\texttt{MLA}})$ are the following summary statistics:
@@
$\widehat{\bm{\theta}}_{\texttt{ss}}=(\frac{1}{n}\sum\limits_{i=1}^{n}y_{i},\frac{1}{n}\sum\limits_{i=1}^{n}\widehat{f}_{i},\frac{1}{N}\sum\limits_{i=n+1}^{n+N}\widehat{f}_{i})\text{ and }\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\texttt{ss}})=\begin{bmatrix}\widehat{\textnormal{Var}}_{l}[Y]/n&\widehat{\textnormal{Cov}}_{l}[Y,\widehat{f}]/n&0\\ \widehat{\textnormal{Cov}}_{l}[Y,\widehat{f}]/n&\widehat{\textnormal{Var}}_{l}[\widehat{f}]/n&0\\ 0&0&\widehat{\textnormal{Var}}_{u}[\widehat{f}]/N\\ \end{bmatrix}$
@@
Moreover, they can be easily obtained by applying the same algorithm $\mathcal{A}$ (mean estimation) to
@@
  • labeled data with observed outcome $\mathcal{A}(Y_{\cal L})\rightarrow[\widehat{\theta}_{{\cal L}},\widehat{\operatorname{Var}}(\widehat{\theta}_{{\cal L}})]=[\frac{1}{n}\sum\limits_{i=1}^{n}y_{i},\widehat{\textnormal{Var}}_{l}[Y]/n]$
@@
  • labeled data with predicted outcome $\mathcal{A}(\widehat{f}_{\cal L})\rightarrow[\widehat{\eta}_{{\cal L}},\widehat{\operatorname{Var}}(\widehat{\eta}_{{\cal L}})]=[\frac{1}{n}\sum\limits_{i=1}^{n}\widehat{f}_{i},\widehat{\textnormal{Var}}_{l}[\widehat{f}]/n]$
[…]
@@
  • bootstrap of labeled data $\mathcal{A}[(Y_{\cal L},\widehat{f}_{\cal L})_{q},q=1,\dots,Q]$ for estimation of $\widehat{\operatorname{Cov}}(\widehat{\theta}_{{\cal L}},\widehat{\eta}_{{\cal L}})=\widehat{\textnormal{Cov}}_{l}[Y,\widehat{f}]/n$. Here, $(Y_{\cal L},\widehat{f}_{\cal L})_{q}$ represents the $q$-th bootstrap of labeled data.
@@
Combining these summary statistics for one-step debiasing $\widehat{\omega}_{0}\widehat{\eta}_{{\cal U}}-(\widehat{\omega}_{0}\widehat{\eta}_{{\cal L}}-\widehat{\theta}_{{\cal L}})$ recovers $\widehat{\theta}_{\texttt{MLA}}$.
@@
To summarize, an algorithm for mean estimation, coupled with resampling, is sufficient for ML-assisted mean estimation. This observation inspired us to generalize this protocol for a broad range of tasks. Our protocol illustrated in fig. 1 only requires three steps: 1) using a pre-trained ML model to predict outcomes for labeled and unlabeled data, 2) applying existing analysis routines to generate summary statistics, and 3) using these statistics in a debiasing procedure to produce statistically valid results in ML-assisted inference.
Figure 1: Workflow of PSPS for Task-Agnostic ML-Assisted Inference.